Run Planning Discussion 2 July, 2015¶
|estimated time||10pm - 6am|
|Lead shifter||Mariyan Bogomilov|
|Shifter 2||Dan Kaplan|
|Physics Shifter||F. Drielsma|
- The priority is placed on the magnet preparation activities.
- Our search and seal procedures represent an impediment to free access of the hall so data collection is suspended until either the case is made for more data prior to the beginning of user run 01b.
- Cooling SSD and SSD is "complete".
- TOF1 will be moved out of the beam line for required Tracker repairs Friday morning. As a result there will be no data collected for the remainder of the user run.
- New shifters Mariyan and Dan will start today. Their primary responsibilities will be monitoring the magnets in the hall overnight.
- see issue #1696
Beam line Optimization¶
- There is evidence that the beam setting corresponding to the run control tag "6-200+MatchDown3_With_DS" produces an increased number of TOF triggers.
- There is a request to retake these data next time beam is available.
- Can the normalization be done with the existing data?
- This will need to wait run 01b for further evaluation.
For "Magnet" shifters¶
Please put your records of the shift including any parameters of interest (such as pressures) on the e-log as with any other shift. If you are writing it down it should be in a place of public record that won't get lost.
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
Access Code: 2155402